OpenStudy (anonymous):

solve the equation for solutions in the interval [0,2pi] 8 sin x-8=-4 help mee plzzzzzzzzzzzzzzzzzzzzzzzzzzz

OpenStudy (amoodarya):

8 sin x-8=-4 8 sin x=8-4 8 sin x=4 sinx=1/2 sinx=sin(30) so x=30+2kpi or x=180-30+2kpi